The percentage of women in the 20-24 age group who have never married went from 55% in 1970 to 73% in 2000 (census Bureau, www.census.gov). Let 1970 be year 0 and 2000 be year 30.
a) Find and interpret the slope of the line through the points (0,55) and (30,73).
b) Find the equation of the line in part (a).
c) What is the meaning of the y-intercept?
d)Use the equation to predict the percentage in 2010.
e) If this trend continues, then in what year will the percentage of women in the 20-24 age group who have never married reach 100%?

Answer by stanbon(75887)

The percentage of women in the 20-24 age group who have never married went from 55% in 1970 to 73% in 2000 (census Bureau, www.census.gov). Let 1970 be year 0 and 2000 be year 30.
a) Find and interpret the slope of the line through the points (0,55) and (30,73).
slope = (73-55)/(30-0) = 0.6
--------------------------------
b) Find the equation of the line in part (a).
% = 0.6(years after 1970)+55
--------------------------------
c) What is the meaning of the y-intercept?
In 1970 the percentage was 55
--------------------------------
d)Use the equation to predict the percentage in 2010.
% = 0.6*40+55
% = 24+55 = 79
---------------------------
e) If this trend continues, then in what year will the percentage of women in the 20-24 age group who have never married reach 100%?
100 = 0.6x+55
45 = 0.6x
x = 75
!00% in 1970+75 = 2045
====================
Cheers,
Stan H.